WARNING: This web is not used anymore. Please use PDBService.MinuteS24Aug06 instead!

Minutes 24 August 06

  • Phone:
    • TRIUMF: Denise
    • RAL: Gordon
    • BNL Iris
    • GridKA: Andreas
    • ATLAS: Sasha
    • Oxford: Stefan
    • CNAF: Barbara

  • CERN:
    • 3D: Dirk, Eva, Zbigniew
    • LHCB: Marco
    • LFC: Jean-Philippe
    • ARDA: Birger

  • Round table site status:
    • Taiwan: db cluster expected to be available end next week.
    • BNL (Iris): problem of the disk space missing has been solved. OEM agents still pending, problem reported to Oracle.
    • CERN (Eva): archivelog mode enabled on the integration RACs.
    • GridKA (Andreas): shared pool memory problem disappeared after database restart.
    • IN2P3: Streams setup for the LHCb experiment ongoing.
    • RAL (Gordon): trhoughput tests re-run: response time passed.
    • CNAF (Barbara): second cluster expected to be available begining next week. Setting up storage (not the definitive one, waiting for the Fiber Channel Storage - expected end of September).
    • NDGF: hw request ongoing.
    • SARA: setting up a test system.
    • PIC: looking for dba contact. Hw almost ready.
    • TRIUMF (Denise): hw ordered, test installation.

  • LFC replication test status:
    • Simulated 40 parallel clients.
    • Rate: 3700 entries per minute. Observed: slave slower than the master.
    • Rate foreseen: 2 or 3 entries per second.
    • Objective for the next steps: achive the rate on which replication is stable (no flow control).

  • LHCB COOL tests:
    • Creation of schema done
    • Replication to RAL correct
    • Replication to GridKA: shared memory problem found and Apply process aborted. After solving the memory problem, apply process was re-enabled, and replication finished without problem. From dba point of view, everything has been replicated. Marco will check from the application point of view.
    • IN2P3: needed to import the contents already replicated to RAL and GridKA.
    • New account to be replicated. Eva will contact the sites and will add it to the replication setup.

  • ATLAS COOL tests:
    • Florbela and Gancho are on holidays until September.
    • They will do the Streams setup.

  • Recovery tests with Streams:
    • Test environment + Streams setup is ready to start the tests.
    • Population will run during some days. We will start with the recover of a destination site: RAL. Schedule to Wednesday.

  • Oracle BUG:
    • Problem has been identified.
    • Workaround available.
    • Patch to be released and applied during the next days!!
Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2006-08-25 - EvaDafonte
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback